# Inflationary Models with Gauss–Bonnet Coupling in Light of ACT Observations

### v1

@@ Abstract
Recent analyses combining Atacama Cosmology Telescope (ACT) data with other cosmological datasets report a higher scalar spectral index $n_{s}$, creating tension with a wide range of inflationary models. Since a Gauss–Bonnet term with a coupling function $\xi(\phi)=3\lambda/[4V(\phi)]$ leaves $n_{s}$ nearly unchanged (up to a field rescaling) while reducing the tensor-to-scalar ratio $r$ by a factor $(1-\lambda)$, so choosing $(1-\lambda)$ sufficiently small effectively removes $r$ as a limiting observable, making it easier for inflationary models to satisfy the latest observational constraints and alleviating this tension. Applying this mechanism to chaotic inflation, E-models, T-models, and hilltop inflation, we find that broad regions of parameter space become consistent with the latest ACT-based CMB constraints. These results demonstrate that Gauss–Bonnet couplings can help bring a broad class of inflationary models into agreement with current CMB measurements.
@@ 1 Introduction
Inflation has become a cornerstone of modern cosmology, providing a compelling mechanism to solve the flatness, horizon, and monopole problems of the standard Big Bang model, while simultaneously generating the primordial perturbations that seed large-scale structures and leave imprints as anisotropies in the cosmic microwave background (CMB) [1, 2, 3, 4]. These primordial perturbations are commonly characterized by two key observables: the scalar spectral index $n_{s}$, which quantifies the scale dependence of scalar perturbations, and the tensor-to-scalar ratio $r$, which measures the relative amplitude of primordial gravitational waves. For a given inflationary potential, both observables are typically expressed in terms of the number of $e$-folds NN between horizon exit and the end of inflation, allowing precise theoretical predictions to be directly compared with observations. A well-known class is the so-called universal attractors [5], for which the scalar spectral index takes the form $n_{s}=1-2/N$. This prediction arises in many models, including E- and T-models [6, 7], $R^{2}$ inflation [1], and Higgs inflation with strong nonminimal coupling [8, 9]. For $N=60$, the universal attractor yields $n_{s}=0.9667$, which lies in excellent agreement with the Planck 2018 result $n_{s}=0.9649\pm 0.0042$ [10].
@@
However, recent observations from the Atacama Cosmology Telescope (ACT) [11, 12], when combined with other datasets , report a higher value of $n_{s}$ compared to Planck alone. A joint analysis of ACT and Planck data (P-ACT) yields $n_{s}=0.9709\pm 0.0038$, while including CMB lensing and Dark Energy Spectroscopic Instrument (DESI) BAO measurements [13, 14] (P-ACT-LB) further increases this to $n_{s}=0.9743\pm 0.0034$ [11, 12]. These latest observational data disfavor the universal attractors at approximately the $2\sigma$ level, thereby creating significant tension for a broad class of inflationary models that predict them. Several approaches have been proposed to address the tension between the latest observational data and inflationary models. These include relaxing the strong-coupling limit in nonminimally coupled inflationary models of the form $\xi f(\phi)R$ [15, 16], incorporating reheating dynamics [17, 18, 19, 20, 21, 22, 23, 24, 25, 26], and investigating alternative inflationary frameworks [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. For a recent overview, see Ref. [50].
@@
In this work, we propose an alternative resolution based on a Gauss–Bonnet term coupled to the inflation field via $\xi(\phi)=3\lambda/[4V(\phi)]$ [51, 52]. This coupling leaves the scalar spectral index $n_{s}$ identical to that of the corresponding canonical model with the same potential under a rescaled field, while suppressing the tensor-to-scalar ratio $r$ by a factor of $(1-\lambda)$. As a result, $r$ can be made sufficiently small without significantly affecting $n_{s}$, effectively removing the tensor-to-scalar ratio as a constraint in model selection. The comparison with observational data can then focus solely on $n_{s}$, substantially relaxing observational tensions and enabling a wide range of inflationary models to remain compatible with the recent P-ACT-LB measurements.
@@
We apply this mechanism to several representative inflationary models, including chaotic inflation, E- and T-models, and hilltop inflation. For each model, we compute the predictions for $n_{s}$ and $r$, determine the allowed parameter space, and compare with the latest observational constraints. Our results indicate that the Gauss–Bonnet coupling can help alleviate the tension between inflationary models and current CMB constraints. Related studies on reconciling inflationary potentials with the P-ACT-LB data using a Gauss–Bonnet term involving specific choices for the coupling function, the potential, or both can be found in Ref. [47], which also incorporates the GW170817 constraints, and in Refs.[48, 49], where the coupling function is taken to have a hyperbolic or exponential form.
@@ 2.1 The background
$S=\frac{1}{2}\int\sqrt{-g}d^{4}x\left[R-g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-2V(\phi)-\xi(\phi)R^{2}_{GB}\right],$ (2.1)
@@
$6H^{2}=\dot{\phi}^{2}+2V+24\dot{\xi}H^{3},$ (2.2)
$2\dot{H}=-\dot{\phi}^{2}+4\ddot{\xi}H^{2}+4\dot{\xi}H(2\dot{H}-H^{2}),$ (2.3)
$\ddot{\phi}+3H\dot{\phi}+V_{,\phi}+12\xi_{,\phi}H^{2}(\dot{H}+H^{2})=0,$ (2.4)
where a dot denotes a derivative with respect to cosmic time $t$, e.g., $\dot{\phi}=d\phi/dt$, and a subscript comma represents an ordinary derivative with respect to the scalar field, e.g., $V_{,\phi}\equiv dV/d\phi$.
@@
$\dot{\phi}\ll V(\phi),\quad|\ddot{\phi}|\ll 3H|\dot{\phi}|,\quad 4H|\dot{\xi}|\ll 1,\quad|\ddot{\xi}|\ll H|\dot{\xi}|.$ (2.5)
To quantify these conditions, we define the Hubble flow parameters $\epsilon_{i}$ and the coupling function flow parameters $\delta_{i}$ as [53, 54]
$\epsilon_{1}=-\frac{\dot{H}}{H^{2}},\quad\epsilon_{i+1}=\frac{d\ln|\epsilon_{i}|}{d\ln a},\quad i\geq 1,$ (2.6)
$\delta_{1}=4\dot{\xi}H,\quad\delta_{i+1}=\frac{d\ln|\delta_{i}|}{d\ln a},\quad i\geq 1.$ (2.7)
Using these slow-roll parameters, the slow-roll conditions (2.5) can be expressed as
$\epsilon_{1}\ll 1,\quad|\epsilon_{2}|\ll 1,\quad|\delta_{1}|\ll 1,\quad|\delta_{2}|\ll 1.$ (2.8)
In this regime, the background equations (2.2)- (2.4) simplify to
$H^{2}\approx\frac{1}{3}V,$ (2.9)
$\dot{H}\approx-\frac{1}{2}\dot{\phi}^{2}-2\dot{\xi}H^{3},$ (2.10)
$\dot{\phi}\approx-\frac{1}{3H}(V_{,\phi}+12\xi_{,\phi}H^{4}).$ (2.11)
@@
$\epsilon_{1}\approx\frac{Q}{2}\frac{V_{,\phi}}{V},\quad\epsilon_{2}\approx-Q\left(\frac{V_{,\phi\phi}}{V_{,\phi}}-\frac{V_{,\phi}}{V}+\frac{Q_{,\phi}}{Q}\right),$ (2.12)
$\delta_{1}\approx-\frac{4}{3}\xi_{,\phi}QV,\quad\delta_{2}\approx-Q\left(\frac{\xi_{,\phi\phi}}{\xi_{,\phi}}+\frac{V_{,\phi}}{V}+\frac{Q_{,\phi}}{Q}\right),$ (2.13)
@@ 2.2 The perturbation
The Mukhanov-Sasaki equation for the scalar perturbation is [55, 56, 57, 58, 59, 60]
$v_{k}^{\prime\prime}+\left(c_{s}^{2}k^{2}-\frac{z_{s}^{\prime\prime}}{z_{s}}\right)v_{k}=0,$ (2.17)
@@
$c_{s}^{2}=1-\Delta^{2}\frac{2\epsilon_{1}+\frac{1}{2}\delta_{1}(1-5\epsilon_{1}-\delta_{2})}{F},$ (2.18)
$z_{s}^{2}=a^{2}\frac{F}{(1-\frac{1}{2}\Delta)^{2}},$ (2.19)
@@
$\Delta=\delta_{1}/(1-\delta_{1}),\quad F=2\epsilon_{1}-\delta_{1}(1+\epsilon_{1}-\delta_{2})+3\Delta\delta_{1}/2.$ (2.20)
@@
$\frac{z_{s}^{\prime\prime}}{z_{s}}=\frac{1}{\tau^{2}}\left(\nu-\frac{1}{4}\right),$ (2.21)
@@
$\nu=\frac{3}{2}+\epsilon_{1}+\frac{2\epsilon_{1}\epsilon_{2}-\delta_{1}\delta_{2}}{4\epsilon_{1}-2\delta_{1}}.$ (2.22)
@@
$\mathcal{P_{R}}$ $=\frac{k^{3}}{2\pi^{2}}\left|\frac{v_{k}}{z}\right|^{2}$ (2.23)
$=2^{2\nu-3}\left[\frac{\Gamma(\nu)}{\Gamma(3/2)}\right]^{2}\frac{(1-\Delta/2)^{2}}{Fc_{s}^{3}}\left(\frac{H}{2\pi}\right)^{2}(1-\epsilon_{1})^{2\nu-1}\left(\frac{c_{s}k}{aH}\right)^{3-2\nu}\left.\right|_{c_{s}k=aH},$
where $\Gamma(x)$ denotes the Gamma function. The scale spectral index is [54]
$n_{s}-1=\frac{d\ln\mathcal{P_{R}}}{d\ln k}=-2\epsilon_{1}-\frac{2\epsilon_{1}\epsilon_{2}-\delta_{1}\delta_{2}}{2\epsilon_{1}-\delta_{1}}.$ (2.24)
Similarly, for tensor perturbations, the Mukhanov-Sasaki equation is [57, 58, 59, 60]
$\frac{d^{2}u^{b}_{k}}{d\tau^{2}}+\left(c_{T}^{2}k^{2}-\frac{z_{T}^{\prime\prime}}{z_{T}}\right)u^{b}_{k}=0,$ (2.25)
@@
$z_{T}^{2}=a^{2}(1-\delta_{1}),\quad c_{T}^{2}=1+\Delta(1-\epsilon_{1}-\delta_{2}).$ (2.26)
@@
$\mathcal{P_{T}}$ $=\frac{k^{3}}{2\pi^{2}}\sum_{b=+,\times}\left|\frac{2u_{k}^{b}}{z_{T}}\right|^{2}$ (2.27)
$=\frac{2^{2\mu}}{(1-\delta_{1})c_{T}^{3}}\left[\frac{\Gamma(\mu)}{\Gamma(3/2)}\right]^{2}\left(\frac{H}{2\pi}\right)^{2}(1-\epsilon_{1})^{2\mu-1}\left(\frac{c_{T}k}{aH}\right)^{3-2\mu}\left.\right|_{c_{T}k=aH},$
with $\mu=3/2+\epsilon_{1}$. The tensor-to-scalar ratio is [54]
$r=\frac{\mathcal{P_{T}}}{\mathcal{P_{R}}}=16\epsilon_{1}-8\delta_{1},$ (2.28)
and the tensor spectral index
$n_{T}=\frac{d\ln\mathcal{P_{T}}}{d\ln k}=-2\epsilon_{1}.$ (2.29)

### v2

@@ Abstract
Recent analyses combining Atacama Cosmology Telescope (ACT) data with other cosmological datasets report a higher scalar spectral index $n_{s}$, creating tension with a wide range of inflationary models. Since a Gauss–Bonnet term with a coupling function $\xi(\phi)=3\lambda/[4V(\phi)]$ leaves $n_{s}$ nearly unchanged (up to a field rescaling) while reducing the tensor-to-scalar ratio $r$ by a factor $(1-\lambda)$, so choosing $(1-\lambda)$ sufficiently small effectively removes $r$ as a limiting observable, making it easier for inflationary models to satisfy the latest observational constraints and alleviating this tension. Applying this mechanism to chaotic inflation, E-models, T-models, and hilltop inflation, we find that broad regions of parameter space become consistent with the latest ACT-based cosmic microwave background (CMB) constraints. These results demonstrate that Gauss–Bonnet couplings can help bring a broad class of inflationary models into agreement with current CMB measurements.
journal: Eur. Phys. J. C
@@ 1 Introduction
Inflation has become a cornerstone of modern cosmology, providing a compelling mechanism to solve the flatness, horizon, and monopole problems of the standard Big Bang model, while simultaneously generating the primordial perturbations that seed large-scale structures and leave imprints as anisotropies in the cosmic microwave background (CMB) Starobinsky:1980te ; Guth:1980zm ; Linde:1981mu ; Albrecht:1982wi . These primordial perturbations are commonly characterized by two key observables: the scalar spectral index $n_{s}$, which quantifies the scale dependence of scalar perturbations, and the tensor-to-scalar ratio $r$, which measures the relative amplitude of primordial gravitational waves. For a given inflationary potential, both observables are typically expressed in terms of the number of $e$-folds NN between horizon exit and the end of inflation, allowing precise theoretical predictions to be directly compared with observations. A well-known class is the so-called universal attractors Kallosh:2013tua , for which the scalar spectral index takes the form $n_{s}=1-2/N$. This prediction arises in many models, including E- and T-models Kallosh:2013hoa ; Kallosh:2013maa , $R^{2}$ inflation Starobinsky:1980te , and Higgs inflation with strong nonminimal coupling Kaiser:1994vs ; Bezrukov:2007ep . For $N=60$, the universal attractor yields $n_{s}=0.9667$, which lies in excellent agreement with the Planck 2018 result $n_{s}=0.9649\pm 0.0042$ Planck:2018jri .
@@
However, recent observations from the Atacama Cosmology Telescope (ACT) ACT:2025fju ; ACT:2025tim , when combined with other datasets , report a higher value of $n_{s}$ compared to Planck alone. A joint analysis of ACT and Planck data (denoted as P-ACT) yields $n_{s}=0.9709\pm 0.0038$, while including CMB lensing and Dark Energy Spectroscopic Instrument (DESI) baryon acoustic oscillation (BAO) measurements DESI:2024uvr ; DESI:2024mwx (denoted as P-ACT-LB) further increases this to $n_{s}=0.9743\pm 0.0034$ ACT:2025fju ; ACT:2025tim . These latest observational data disfavor the universal attractors at approximately the $2\sigma$ level, thereby creating significant tension for a broad class of inflationary models that predict them. Several approaches have been proposed to address the tension between the latest observational data and inflationary models. These include relaxing the strong-coupling limit in nonminimally coupled inflationary models of the form $\xi f(\phi)R$ Kallosh:2025rni ; Gao:2025onc , incorporating reheating dynamics Liu:2025qca ; Haque:2025uis ; Zharov:2025evb ; Haque:2025uri ; Drees:2025ngb ; Ballardini:2024ado ; Ye:2025idn ; SidikRisdianto:2025qvk ; Zharov:2025zjg ; Chakraborty:2025oyj , and investigating alternative inflationary frameworks Kim:2025dyi ; Wolf:2025ecy ; He:2025bli ; Gialamas:2025kef ; Frob:2025sfq ; Dioguardi:2025vci ; Brahma:2025dio ; Berera:2025vsu ; Aoki:2025wld ; Dioguardi:2025mpp ; Salvio:2025izr ; Gialamas:2025ofz ; Peng:2025bws ; Pallis:2025nrv ; Katsoulas:2025mcu ; Byrnes:2025kit ; Maity:2025czp ; Addazi:2025qra ; Mondal:2025kur ; Odintsov:2025wai ; Yogesh:2025wak ; Zahoor:2025nuq ; Mohammadi:2025gbu ; Odintsov:2025eiv ; Oikonomou:2025xms ; Hai:2025wvs ; Kuralkar:2025zxr . For a recent overview, see Ref. Kallosh:2025ijd .
@@
In this work, we propose an alternative resolution based on a Gauss–Bonnet term coupled to the inflation field via $\xi(\phi)=3\lambda/[4V(\phi)]$ Guo:2009uk ; Jiang:2013gza . This coupling leaves the scalar spectral index $n_{s}$ identical to that of the corresponding canonical model with the same potential under a rescaled field, while suppressing the tensor-to-scalar ratio $r$ by a factor of $(1-\lambda)$. As a result, $r$ can be made sufficiently small without significantly affecting $n_{s}$, effectively removing the tensor-to-scalar ratio as a constraint in model selection. The comparison with observational data can then focus solely on $n_{s}$, substantially relaxing observational tensions and enabling a wide range of inflationary models to remain compatible with the recent P-ACT-LB measurements. For additional recent work on the Gauss–Bonnet couplings, see e.g. Refs. Yi:2018gse ; Odintsov:2025kyw ; Odintsov:2023aaw ; Odintsov:2020mkz ; Oikonomou:2024etl ; Nojiri:2017ncd ; Gangopadhyay:2022vgh ; Khan:2022odn ; Yogesh:2025hll ; Yogesh:2024zwi ; Yogesh:2024mpa ; Luo:2024vls ; Hussain:2025vbo .
@@
We apply this mechanism to several representative inflationary models, including chaotic inflation, E- and T-models, and hilltop inflation. For each model, we compute the predictions for $n_{s}$ and $r$, determine the allowed parameter space, and compare with the latest observational constraints. Our results indicate that the Gauss–Bonnet coupling can help alleviate the tension between inflationary models and current CMB constraints. Related studies on reconciling inflationary potentials with the P-ACT-LB data using a Gauss–Bonnet term involving specific choices for the coupling function, the potential, or both can be found in Ref. Odintsov:2025wai , which also incorporates the GW170817 constraints, and in Refs.Yogesh:2025wak ; Zahoor:2025nuq , where the coupling function is taken to have a hyperbolic or exponential form.
@@ 2.1 The background
$S=\frac{1}{2}\int\sqrt{-g}d^{4}x\bigg[R-g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-2V(\phi)-\xi(\phi)R^{2}_{GB}\bigg],$ (1)
@@
$6H^{2}=\dot{\phi}^{2}+2V+24\dot{\xi}H^{3},$ (2)
$2\dot{H}=-\dot{\phi}^{2}+4\ddot{\xi}H^{2}+4\dot{\xi}H(2\dot{H}-H^{2}),$ (3)
$\ddot{\phi}+3H\dot{\phi}+V_{,\phi}+12\xi_{,\phi}H^{2}(\dot{H}+H^{2})=0,$ (4)
where $H=\dot{a}/a$ is the Hubble parameter with a(t) being the cosmic scale factor. A dot denotes a derivative with respect to cosmic time $t$, e.g., $\dot{\phi}=d\phi/dt$, and a subscript comma represents an ordinary derivative with respect to the scalar field, e.g., $V_{,\phi}\equiv dV/d\phi$.
@@
$\dot{\phi}\ll V(\phi),\;|\ddot{\phi}|\ll 3H|\dot{\phi}|,\;4H|\dot{\xi}|\ll 1,\;|\ddot{\xi}|\ll H|\dot{\xi}|.$ (5)
To quantify these conditions, we define the Hubble flow parameters $\epsilon_{i}$ and the coupling function flow parameters $\delta_{i}$ as Schwarz:2001vv ; Guo:2010jr
$\epsilon_{1}=-\frac{\dot{H}}{H^{2}},\quad\epsilon_{i+1}=\frac{d\ln|\epsilon_{i}|}{d\ln a},\quad i\geq 1,$ (6)
$\delta_{1}=4\dot{\xi}H,\quad\delta_{i+1}=\frac{d\ln|\delta_{i}|}{d\ln a},\quad i\geq 1.$ (7)
Using these slow-roll parameters, the slow-roll conditions (5) can be expressed as
$\epsilon_{1}\ll 1,\quad|\epsilon_{2}|\ll 1,\quad|\delta_{1}|\ll 1,\quad|\delta_{2}|\ll 1.$ (8)
In this regime, the background equations (2)- (4) simplify to
$H^{2}\approx\frac{1}{3}V,$ (9)
$\dot{H}\approx-\frac{1}{2}\dot{\phi}^{2}-2\dot{\xi}H^{3},$ (10)
$\dot{\phi}\approx-\frac{1}{3H}(V_{,\phi}+12\xi_{,\phi}H^{4}).$ (11)
@@
$\epsilon_{1}\approx\frac{Q}{2}\frac{V_{,\phi}}{V},\quad\epsilon_{2}\approx-Q\left(\frac{V_{,\phi\phi}}{V_{,\phi}}-\frac{V_{,\phi}}{V}+\frac{Q_{,\phi}}{Q}\right),$ (12)
$\delta_{1}\approx-\frac{4}{3}\xi_{,\phi}QV,\quad\delta_{2}\approx-Q\left(\frac{\xi_{,\phi\phi}}{\xi_{,\phi}}+\frac{V_{,\phi}}{V}+\frac{Q_{,\phi}}{Q}\right),$ (13)
@@ 2.2 The perturbation
The Mukhanov-Sasaki equation for the scalar perturbation is Mukhanov:1985rz ; Sasaki:1986hm ; Hwang:1999gf ; Cartier:2001is ; Hwang:2005hb ; DeFelice:2011zh
$v_{k}^{\prime\prime}+\left(c_{s}^{2}k^{2}-\frac{z_{s}^{\prime\prime}}{z_{s}}\right)v_{k}=0,$ (16)
@@
$c_{s}^{2}=1-\Delta^{2}\frac{2\epsilon_{1}+\frac{1}{2}\delta_{1}(1-5\epsilon_{1}-\delta_{2})}{F},$ (17)
$z_{s}^{2}=a^{2}\frac{F}{(1-\frac{1}{2}\Delta)^{2}},$ (18)
@@
$\Delta$ $=\delta_{1}/(1-\delta_{1}),$ (19)
$F$ $=2\epsilon_{1}-\delta_{1}(1+\epsilon_{1}-\delta_{2})+3\Delta\delta_{1}/2.$ (20)
@@
$\frac{z_{s}^{\prime\prime}}{z_{s}}=\frac{1}{\tau^{2}}\left(\nu-\frac{1}{4}\right),$ (21)
@@
$\nu=\frac{3}{2}+\epsilon_{1}+\frac{2\epsilon_{1}\epsilon_{2}-\delta_{1}\delta_{2}}{4\epsilon_{1}-2\delta_{1}}.$ (22)
@@
$\mathcal{P_{R}}$ $=\frac{k^{3}}{2\pi^{2}}\left|\frac{v_{k}}{z}\right|^{2}$ (23)
$=2^{2\nu-3}\left[\frac{\Gamma(\nu)}{\Gamma(3/2)}\right]^{2}(1-\epsilon_{1})^{2\nu-1}$
$\left.\times\frac{(1-\Delta/2)^{2}}{Fc_{s}^{3}}\left(\frac{H}{2\pi}\right)^{2}\left(\frac{c_{s}k}{aH}\right)^{3-2\nu}\right|_{c_{s}k=aH},$
where $\Gamma(x)$ denotes the Gamma function. The scale spectral index is Guo:2010jr
$n_{s}-1=\frac{d\ln\mathcal{P_{R}}}{d\ln k}=-2\epsilon_{1}-\frac{2\epsilon_{1}\epsilon_{2}-\delta_{1}\delta_{2}}{2\epsilon_{1}-\delta_{1}}.$ (24)
Similarly, for tensor perturbations, the Mukhanov-Sasaki equation is Hwang:1999gf ; Cartier:2001is ; Hwang:2005hb ; DeFelice:2011zh
$\frac{d^{2}u^{b}_{k}}{d\tau^{2}}+\left(c_{T}^{2}k^{2}-\frac{z_{T}^{\prime\prime}}{z_{T}}\right)u^{b}_{k}=0,$ (25)
@@
$z_{T}^{2}=a^{2}(1-\delta_{1}),\quad c_{T}^{2}=1+\Delta(1-\epsilon_{1}-\delta_{2}).$ (26)
@@
$\mathcal{P_{T}}$ $=\frac{k^{3}}{2\pi^{2}}\sum_{b=+,\times}\left|\frac{2u_{k}^{b}}{z_{T}}\right|^{2}$ (27)
$=\frac{2^{2\mu}}{(1-\delta_{1})c_{T}^{3}}\left[\frac{\Gamma(\mu)}{\Gamma(3/2)}\right]^{2}$
$\times\left.(1-\epsilon_{1})^{2\mu-1}\left(\frac{H}{2\pi}\right)^{2}\left(\frac{c_{T}k}{aH}\right)^{3-2\mu}\right|_{c_{T}k=aH},$
with $\mu=3/2+\epsilon_{1}$. The tensor-to-scalar ratio is Guo:2010jr
$r=\frac{\mathcal{P_{T}}}{\mathcal{P_{R}}}=16\epsilon_{1}-8\delta_{1},$ (28)
and the tensor spectral index is
$n_{T}=\frac{d\ln\mathcal{P_{T}}}{d\ln k}=-2\epsilon_{1}.$ (29)